Neptune Corporation owns 70 percent of Pluto Company's stock. On July 1, 20X4, Neptune sold a piece of equipment to Pluto for $56,350. Neptune had purchased this equipment on January 1, 20X1, for $63,000. The equipment's original 15-year estimated total economic life remains unchanged. Both companies use straight-line depreciation. The equipment's residual value is considered negligible.

1. Based on the information provided, the gain on the sale of equipment eliminated in the consolidated financial statements for 20X4 is

A. $5950

B. $8050

C. $10150

D. $14700

Based on the information provided, while preparing the 20X4 condolidated income statement, depreciation expense will be:

A. credited for $350.

B. debited for $350.

C. credited for $700

D. debited for $700.
Answer #1

1. Option B $ 8050

Equipment Cost as on 01.jan. 2014= 63000$

Useful life = 15 years

SLM Depreciation Per Annum =( cost of Machine-Residual Value)/Useful Life

= (63000$-0)/15= 4200$

Total depreciation for the period from 1.1.2011 to 1.7.2014 is 14700$

W.D.V of Equipment as on 1.july.2014=cost of equipment-Total Depreciation for the period

=63000-14700= 48300$

Gain on Equipment= sold price -W.D.V as on 1.july.2017

56350-48300= 8050$

2) option D. Debited for $700

Gain On equipment= 8050$

Balance period 11.5 years

Annual depreciation debited value to Income statement is 700$( 8050/11.5 year)